10) A 6.0ohms and a 12ohms resistor are connected in parallel to a 36-V battery. What power is dissipated by the 6.0ohms resistor?
A) 220 W
B) 48 W
C) 490 W
D) 24 W
Answer: A
explanation
Answer #1

Power throgh 6 ohm = V^2/R

V =36 V as it is parallel connection

So

P = 36^2/6 = 216 W (ans)

So correct option is A)220 W (approx) (ans)